### v2

# Vertex Imaging Hadron Calorimetry Using AI/ML Tools

Abstract
The fluctuations in energy loss to processes that do not generate measurable signals, such as binding energy losses, set the limit on achievable hadronic energy resolution in traditional energy reconstruction techniques. The correlation between the number of hadronic interaction vertices in a shower and invisible energy is found to be strong and is used to estimate invisible energy fraction in highly granular calorimeters in short time intervals (<10 ns). We simulated images of hadronic showers using GEANT4 and deployed a neural network to analyze the images for energy regression. The neural network-based approach results in significant improvement in energy resolution, from 13% to 4% in the case of a Cherenkov calorimeter for 100 GeV pion showers. We discuss the significance of the phenomena responsible for this improvement and the plans for experimental verification of these results and further development.
1 Introduction
Main goal of hadron calorimeter is to measure high energy quarks and gluons produced in particle collisions in experiments. Each quark or gluon appears in detector as a stream of hadrons, called jet, which primarily consists of photons from $\pi^{0}$ decays and other hadrons, $\pi^{\pm},K^{\pm,0},n,p$. The former forms the electromagnetic shower (em) and later forms the hadronic shower (had) in a calorimeter. The calorimeter response to a photon or a hadron tend to be different, typically (photon response)$>$(hadron response). We discuss the role of “invisible energy” in response differences in the next sections. Two methods have been introduced to overcome the invisible energy problem. TheCompensation method was introduce in the late 1970s and implemented the average energy compensation in hardware. The Dual Readout (DR) method was introduced in early 2000s and apply event-by-event compensation using scintillation signal and Cherenkov signal in offline software.
We started the CaloX project at the Advanced Particle Detector Laboratory at TTU in 2019. Goal was to develop a novel calorimetry concept which utilizes new technologies -AI/ML techniques, high granularity of calorimeter, fast photo detector and high density electronics. We aimed at the signal collection time less than 10 ns, which is well below the collision time frequency at the LHC (40 MHz, 25 ns) and the future collider FCC-hh. We reported our earlier results in RefJ01 ; RefJ02 . In the following sections, we discuss the sources of invisible energy and our concept of fast calorimetry for the future collider experiments.
2 Vertex Imaging Calorimetry
2.1 Invisible Energy at Vertex
A hadronic shower consists of many hadron-nucleus interactions where the interaction point can be thought of as a vertex with tracks emerging from it (Figure 1). At each vertex, a fraction of energy of the incoming particle is lost to break up nuclei (binding energy losses) and generate many low energy protons and neutrons that are undetectable, either stopped in the absorber, or below the Cherenkov threshold. Ionization calorimeter detects some of those slow neutrons and protons, resulting in smaller invisible energy. Since the binding energy loss is strongly correlated with the number of neutrons and protons, Compensation and DR methods detect fraction of those and use them in estimating the invisible energy.
Figure 1: (a) A sketch of a pion interacting with a nucleus and (b) Image of a vertex and tracks from a proton-nucleus interaction in emulsion (image from RefJ04 ). A 30-GeV proton from right interacts with the emulsion stack. Fast pions and fast spallation protons appear as thin tracks in the forward region and slow protons as thick isotropic tracks. The neutrons are invisible. (c) Break down of source of the invisible energy in Cherenkov calorimeter and (d) the velocity ($\beta$) of mesons, baryons and nuclear fragments in showers in $\pi^{+}Cu$ interaction. The Cherenkov thresholds for the index of refraction 1.49 (1.80) for plastic (sapphire) fibers are shown. Particles below the threshold are not visible in Cherenkov calorimeter. Other sources are neutrinos, muons, mass of baryons produced in baryon productions and leakage of shower due to the finite size of the calorimeter. The average invisible energy is 31% of the beam energy in this example.
2.2 Vertex Counting for Invisible Energy Estimation
There are many vertices in a hadron shower and the invisible energy at each vertex varies. It seems a daunting task to estimate the total invisible energy from these complex hadronic shower images. We simulated ionization and Cherenkov signal using GEANT4 and found a clear correlation between the invisible energy and the number of inelastic hadron interaction vertices (Figure 2). This correlation suggests a possibility to estimate the invisible energy by just counting the number of vertices in the shower. Some AI/ML tools are known to be very efficient to analyze “images.” We used Convolutional Neural (CNN) and Graph Neural Networks (GNN) RefJ03 to estimate the invisible energy from 3D shower images for the beam energy reconstruction.
Figure 2: (a) The correlation between the invisible energy and the number of hadronic vertices within the hadronic shower for 20 and 200 GeV pions where the signal is based on ionization, and (b) for 30, 100, and 200 GeV pions in the Cherenkov signal in copper absorber (Figures from RefJ02 ).
2.3 Energy Reconstruction with GNN
We simulated 3D images of Cherenkov signal in a fiber calorimeter using GEANT4 with the FTFP-BERT physics list. Fibers of 1 mm diameter were placed in 2-m long Cu absober with the separation of fibers 1.5 mm (center-to-center distance). Production points of Cherenkov photons captured in fibers were recorded and fed into a GNN for energy regression. The energy scale was restored and the energy resolution was improved by more than a factor of two (Figure 3). This demonstrates that we can reconstruct hadron energy with good precision using only the fast or Cherenkov component of the shower. Furthermore, the higher index of refraction ($n$=1.80) improves the resolution compared to ($n=$1.49), indicating the importance of detecting slow isotropic protons. This observation is consistent with a result from our previous analysis with CNN Class Activation Mapping in ionization calorimeter, which also indicated the importance of slow protons in transverse direction around vertex RefJ01 . Scintillation calorimeter detects these slow protons. We simulated scintillation fiber calorimeter with same fiber configuration as Cherenkov fibers. The signal integration time was limited to 10 ns. We observed improvement in resolution (Figure 3(c)).
Figure 3: (a) Response and energy resolution of the Cherenkov fiber calorimeter for 4-150 GeV $\pi^{\text{+}}$ with the traditional method (simple sum) and GNN energy regression. Two cases of the index refraction were assumed for the clear “Cherenkov” fiber.
3 Choice of Hadron Calorimeter
3.1 Absorber Material
The binding energy losses in Cu are much smaller compared to that of in other common absorber (e.g. Pb, W, and U). In addition, much fewer neutrons are produced in Cu making it a preferred absober material for imaging calorimetry.
3.2 Sampling and Signal Integration Time
The performances of Cu-fiber calorimeters are compared with other types detectors in Figure 4, and we draw the following conclusions so far from our work: $a$) it is clear that fine sampling is key to good energy resolution as shown in Figure 4; $b$) the short integration time is desirable for the mitigation of beam-induced-background at a future Muon Collider and pile-up effects at high rate hadron colliders, $c$) the experience from the CMS forward calorimeter shows that Cherenkov calorimeter with quartz fibers is radiation-hard, and $d$) as we argue, the use of proper neural networks results in good energy resolution. The scintillation fiber calorimeter with similarly tuned neural network, where the contribution of the slow protons to signal is sizable, provides the best energy resolution, see Figure 4.
Figure 4: The compilation of hadronic energy resolution performance of various operating, planned, or simulated calorimeter at 100 GeV. Note that the longer integration calorimeters are grouped on the top, and the shorter integration times, a key ingredient in our work, are group in the lower part of the plot. (RefJ05 ; RefJ06 ; RefJ07 ; RefJ08 ; RefJ09 )
3.3 Homogeneous Calorimeter
Homogeneous calorimeter is an “ultimate sampling calorimeter.” We simulated Cherenkov signal in a block of PbWO4 crystal ($n=$2.2) and a Cu block, which was treated as an imaginary crystal of $n=$1.49. The energy resolution from Cu with CNN is better than that from PbWO4 in both scintillation and Cherenkov cases (Figure 5). Segmented âCrystalâ of 2$\times$2$\times$2 cm3 does not provide very significant improvement in resolution over the densely packed fiber calorimeter. This indicates that the resolution seen in the fiber calorimeter case was not limited by the sampling and rather it is limited by the method of the invisible energy estimation with CNN/GNN. Further analysis of the effectiveness of CNN/GNN may provide clues to improve the resolution of the fiber calorimeter and reveal the limits of homogeneous hadron calorimetry.
Figure 5: The resolution of of the Cherenkov crystal calorimeter for 4-150 GeV $\pi^{+}$. (a) PbWO4 ($n=$2.2) and (b) "Cu crystal". A Cu block is treated as an imaginary crystal of $n=$1.49.
4 Potential of Cherenkov Fiber Calorimeter
4.1 HG-DREAM
The original DREAM module was constructed in 2002-2003 at TTU and used at CERN to test the DR calorimetry in 2004-2012. It was brought back to TTU in 2023 and will be used as a test bed of new technologies and new ideas. The module has been refurbished with SiPMs for highly granular readout and data taking with cosmic muons has started. The module will be placed in test beams to verify the Vertex Imaging Calorimetry concept and its further development. See Figure 6 for fiber readout configurations and RefJ10 for more about HG-DREAM.
Figure 6: The DREAM module consists of $\sim$5,000 Cu rods of 2-m long and 4$\times$4 mm2 cross section with a 2.5-mm diameter hole. (a) Fiber configurations: In the original DREAM, each PMT received signal from a fiber bundle from 270 rods. In HG-DREAM, 3 (12) rods are read out by a single SiPM at the center (outer) region of the module, respectively. Straight Cherenkov and scintillation fibers are placed in every hole. An alternate configuration using helix fiber is under consideration for TOF measurement.
4.2 TOF measurement
Fiber calorimeter is a 2D ($x,y$) segmented calorimeter. The arrival time of signal at SiPM gives third coordinate ($z$) (see equation in Figure 6). The original scheme of “Longitudinal Segmentation with Timing” assumes $t0$ and $v$ are known, i.e. $t0$ from the beam collision time from the accelerator and $v=c/n$, where $c$ is the speed of light in vacuum and $n$ the index of refraction of the fiber (1.49 for plastic fiber). With a pair of fibers with different velocities $VA$ and $VB$, we may measure both $t0$ and $z$. We simulated helix fibers around straight fibers. Since the path length of light in helix is longer, the effective velocity along the $z$-axis is slower, $VB(z)=VA(z)/1.38$ in the simulated case, where 1.38 is the “helix factor” corresponding to the path length difference. The reconstructed $t0$ is shown Figure 6. This suggests a possibility to build a calorimeter with the TOF measurement resolution up to 10 ps in its full body. Instead of helix fiber, fibers with higher index of refraction, for example Sapphire, $n=$1.78, may give similar effect.
4.3 3D Local Clustering with AI/ML
The Cherenkov signal produces clear and narrow image of hadronic shower comparing to ionization signal (Figure 7). The AI/ML approaches may use these cleaner shower images for 3D pattern recognition of complex event structure, for example:
  • reconstruction of individual particle in jet (then, link to inner tracks for the particle flow reconstruction).
  • analysis of jet substructure to identify boosted $W/Z$ / top / Higgs in multi-TeV collider experiments.
Furthermore, it is conceivable to build an AI-ASIC system on detector to do a series of event reconstruction tasks using the very localized data, for example:
  • pulse train analysis to form hit(x,y,z,E) per channel.
  • clustering of hits to form 3dCluster(x,y,z,E’) per vertex in hadronic shower.
  • linking of 3dClusters to form caloParticle corresponding to incident particle
  • energy regression on each caloParticle
The data from the AI-ASIC system will be caloPraticle(x,y,z,E"). If such system is realized, the transmission of data from the front-end to the back-end will be greatly reduced.
Figure 7: Example of shower images of “pseudo jet.” Three particle hit the calorimeter 10 cm apart. The image with the Cherenkov signal (left) is cleaner than with ionization signal (right). Assignment of the cluster of hits to each incident particle may be performed more effectively with the clean Cherenkov signal rather than the ionization signal.
5 Conclusion
Vertex Imaging Calorimetry using AI/ML may provide very fast hadron calorimetry at the future collider experiments. The integration time of signal will be below 10 ns. The Dual Cherenkov Fiber Readout has a good potential to build a 4D ($xyzt;E$) imaging calorimeter with time resolution of O(10 ps) in its full body. We used simplistic CNN and GNN algorithms thus far. More optimized use of AI/ML will bring further improvement to future calorimetry.